Problem
Before going to school, Eudora ran from her home to a secret laboratory at an average speed of 12 km. Since she was running late, she then took one of her jetpacks and flew to her school at an average speed of 76 km. Eudora traveled a total distance of 120kilometers, and the entire trip took 2 hours.How long did Eudora spend running, and how long did she spend flying using her jetpack?

Answer by jorel555(1290)

Let n represent the time Eudora used her jetpack. Then:
76n+12(2-n)=120
64n=96
n=96/64=1.5 hours with the jetpack, and .5 hours running. ☺☺☺☺